john is twice as old as peter. peter is 5 years older than alice. in 5 years john will be 3 times as old as alice. how old is peter

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
"john is twice as old as peter."
j = 2p
:
" peter is 5 years older than alice."
p = a + 5
or
a = p - 5
:
" in 5 years john will be 3 times as old as alice."
j + 5 = 3(a+5)
j + 5 = 3a + 15
j = 3a + 15 - 5
j = 3a + 10
replace j with 2p
2p = 3a + 10
replace a with (p-5)
2p = 3(p-5) + 10
2p = 3p - 15 + 10
2p - 3p = -5
-p = -5
p = 5 yrs is Peter's age
One problem with this solution however, Alice would be 0 yrs, John is 10